# Domain Specific Pruning of Large Mixture-of-Experts Models with Few-shot Demonstrations

### v1

@@ Appendix A Experiment Details
As shown in the previous analysis in Section 3.2, datasets within the same domain can identify the important experts on other datasets. Thus, for different domains and tasks, we select one dataset as a calibration set, which are shown in Table 6. For mixed-domain pruning, we select the scores calculated on each 25-shot calibration set and average the normalized scores as the final scores of experts. Additionally, all the experiments are conducted in one 8×8$\times$8 × H200 GPU.
@@
Domain Calibration Set
Math AIME 2023
Coding LiveCodeBench-V3
Science GPQA-Main
Agent Dev Set of AgentBench-OS
Appendix B Analysis of Perturbation-based Pruning Method

### v2

@@ Appendix E Experiment Details
As shown in the previous analysis in Section 3.2, datasets within the same domain can identify the important experts on other datasets. Thus, for different domains and tasks, we select one dataset as a calibration set, which is shown in Table 8. For mixed-domain pruning, we choose the scores calculated on each 25-shot calibration set and average the normalized scores as the final scores of experts. For the evaluation of FinanceIQ, we randomly select 1000 samples with the seed of 42 since the test set is too large.
Additionally, all the experiments are conducted in one 8×8$\times$8 × H200 GPU. We set the maximum context length to 32K, the temperature to 0.6, and the top-p sampling value to 0.95 for most benchmarks (temperature as 0.2 for LiveCodeBench). To ensure statistical reliability, most benchmark is evaluated independently 5 times (32 times for math benchmarks), and we report the average performance of pass@1.
@@
Domain Calibration Set License
Math AIME 2023 cc-by-nc-sa-4.0
Coding LiveCodeBench-V3 MIT
Science GPQA-Main MIT
Agent Dev Set of AgentBench-OS Apache-2.0 license
Finance Dev Set of FinanceIQ MIT
Medical Dev Set of USMLE cc-by-nc-sa-4.0
Appendix F Analysis of Perturbation-based Pruning Method
@@
Appendix G Further Experimental Analysis
G.1 Expert Overlap with Different Metrics
We calculate the pairwise overlap of the top-128 experts chosen by various metrics across three datasets: AIME23, GPQA-main, and LiveCodeBench-V3. The results are presented as a heatmap in Figure 9. Our analysis reveals that the experts selected based on gating scores and frequency exhibit a remarkably high overlap, approximately $90\%$. However, the experts identified by our method show significant divergence from these metrics, underscoring the influence of incorporating expert outputs and residual changes in our approach. We posit that the large discrepancy among experts results in a huge performance gap between models.
Figure 9: Overlap of Top-128 Experts Selected via Different Metrics.
G.2 Results of Pruning with Different Ratios at Different Layers
To further investigate the pruning performance of our method, we employed layer-wise dynamic pruning. Specifically, we first normalized the expert scores for each layer and ranked all experts across all layers. Subsequently, we pruned the top experts using different pruning ratios. As shown in Table 9, we observe that employing layer-wise dynamic pruning does not always lead to better performance, and the performance on LiveCodeBench even drops significantly. In addition, the dynamic compression ratio leads to deployment difficulties, as it requires different numbers of experts for each layer. Thus, we suggest just employing a fixed pruning ratio across all layers.
Table 9: Performance comparison of EASY-EP with and without employing layer-wise compression ratios.
Model Ratio Layer AIME2024 GPQA LiveCodeBench
DeepSeek-R1 0.5 $\times$ 79.17 70.12 61.11
0.5 79.33 65.15 44.31
0.25 $\times$ 72.67 67.47 42.51
0.25 65.33 66.16 28.74
DeepSeek-V3-0324 0.5 $\times$ 55.21 65.25 46.71
0.5 58.67 62.63 38.23
0.25 $\times$ 53.12 57.35 27.99
0.25 58.67 60.10 21.56
G.3 Throughput with Different Lengths
To further explore the throughput changes on pruning models, we consider three additional length settings: (1) 1K+4K: 1K input length and 4K output length; (2) 4K+1K: 4K input length and 1K output length; and (3) 4K+4K: 4K input length and 4K output length. Figure 10 (Right) presents results for the other settings. Compared with 1K+1K, either increasing the length of input or output leads to lower throughput and throughput acceleration after pruning. Additionally, in long output scenarios with an equivalent total sequence length, the 128-expert configuration achieves a significantly higher acceleration ratio (2.91$\times$ under the 1K+4K setting, compared to 2.52$\times$ under the 4K+1K setting).
Figure 10: Total throughput across different numbers of experts.